OpenStudy (anonymous):

Write y = 2/3x + 3 in standard form using integers

OpenStudy (anonymous):

\[ y = \frac 2 3 x + 3\\ 3 y = 2 x +9 \] Can you finish it now?

OpenStudy (anonymous):

\[ 2x -3 y + 9=0 \]